Telescope and AGw Commissioning at LFBG with IRTC - 16-17 April 2009 MST

Observers: J. Hill (Tucson), MPedani (@LBT)
Telescope Operator: W.Wack
Telescope Support: J. Rousseau

Summary

This is some bonus observing borrowed from LBC technical observing to test the bug fixes in TCS Build 24, and to try to get some better AGW transformation data. Heavy overcast on Mt. Graham, so no chance of actual observing in twilight. Utility building chiller failed shortly after that.

Jon R. got the chillers semi-recovered by midnight, and the sky improved to only partly cloudy. We opened for about 3 hours, but spent most of the time fighting other problems and changing thermal conditions.

Preparations

TCS Build 24 installed on mtn

Details

Initial Authorization

17:53 pointto M67astrom TelM="ACTIVE" PA=0 gs=34

Apr 16 17:53:11 lbtmu05 irc[25831]: PresetTelescope execution failed: Requested side not allowed for this instrument / focal station combination

17:56 pointto M67astrom TelM="TRACK" PA=0 gs=34 - failed with above message again

Change IIF authorization to LUCIFER

17:57 pointto M67astrom TelM="TRACK" PA=0 gs=34 - succeeded

Change IIF authorization to IRTC

17:59 pointto M67astrom TelM="TRACK" PA=0 gs=34 - failed

18:00 authorize IRTC (from IRAF)

18:01 pointto M67astrom TelM="TRACK" PA=0 gs=34 - succeeded

There must be some funny first time initialization problem going on between IRAF/IRC/IIF. Norm was watching me do this, but neither of us understand it.

Test passing of the guidestar coordinates

18:02 pointto M67astrom TelM="ACTIVE" PA=0 gs=34 - failed because CCD saturated (a common problem in a closed dome with lights on)

Turn down lights in dome,

18:03 pointto M67astrom TelM="ACTIVE" PA=0 gs=35 - of course there are no stars because the dome is closed. But the test is a success because we are able to run Build 24 and talk to the old version of oacserver.

18:05 pointto M67astrom TelM="ACTIVE" PA=0 gs=21

Analyzing the log data later in the evening showed that the coordinates of the guidestar are now getting through as advertised.

Test HBS setpoint in ECS

ECS HBS setpoint is -7.78 degC. Lower C-Ring steel temp is +0.22 degC, so that seems to be working. (still need to optimize the offset value)

Test collimation update/offload after telescope was parked.

18:13 - azslewtohold 135. elslewtohold 63. It isn't obvious that IIF changed the state of anything when Wayne slewed the telescope.

18:14 Remove M2 Offload corrections from SX primary in PSF GUI.

18:05 pointto M67astrom TelM="ACTIVE" PA=0 gs=30 - This applied a little bit of M2 offload, tolerable in this case because the mount was closeby. Bug still exists.

Test updating of M3 selector position

18:20 Send M3 selector to 8.2 degrees. (versus nominal 8.35 degrees for LFBG)

18:22 pointto M67astrom TelM="ACTIVE" PA=0 gs=80 - OSS correctly repositioned the M3 selector.

18:23 Send M3 selector to 8.33 degrees.

18:24 pointto M67astrom TelM="ACTIVE" PA=0 gs=48 - Correctly repositioned the M3 selector.

Test the "show" flag for irc

18:30 authorize telescope as LBC on both sides. Then run the following commands from the IRAF window talking to irc.

IRTC> !irc -show
Open connections:
IRTC

Exit 67
IRTC> authorize LUCIFER
   irc Authorize 'bentGregorianFront left' LUCIFER

   to examine the log: page LOGs/authorize.log
IRTC> !irc -show
Open connections:
IRTC

Exit 67
IRTC> authorize LBC
   irc Authorize 'prime left' LBC

   to examine the log: page LOGs/authorize.log
IRTC> !irc -show
Open connections:
IRTC

Exit 67

Marco running LBC tests with closed dome.

There were a few open minutes with LBC around midnight, but we closed the shutter doors to work on the LBC-Red rotator.

Re-Open

After 3 hours of "coaching" by Jon R, utility chillers are partly working enough to cool the HBS oil. LBC-Red rotator appears stuck. Sky is mostly cloudy. We'll see if we can do anything with IRTC, but it doesn't look hopeful. Right shutter door would not reopen.

00:40 Re-open Left Shutter Door T=-10.8 degC, D=-12.8 degC, but -8.6 degC inside the dome, SX primary at -3.9.

00:45 pointto M5_0460 ACQUIRE gs=0 PAM- - We were defocussed, but not much coma to start with. Wayne added +1 mm Zglobal to get rough focus.

00:51 ptspiral_irtc.csh - Adjust IE=-24 CA=-162 @ AZ=227 EL=45 (encoders were reset yesterday)

01:00 pointto M5_0460 ACTIVE gs=0 PA=0 (enough cluds that we can do active optics on this bright star - guider is complaining that it is too dim) Seeing on the guider is 1.5 - 2 arcsec, plus the mirror has the expected high-order stuff without mirror ventilation.

The zoomed WFS cam view on the GCSGUI is nice.

01:08 pointto M5_0619 ACTIVE gs=0 PA=0 - star not found.

01:13 pointto M5_0619 ACQUIRE gs=0 PAM-

01:15 ptspiral_irtc.csh - found a faint star at IE=-12 CA=-12

01:21 pointto M5_0619 ACTIVE gs=0 PA=0

Testing Azimuth Scale

01:24 pointto M5_0619 ACQUIRE gs=0 PAM- IE=-15 CA=-7 (well-centered, but drifted from 10 minutes ago)

01:28 azslewtohold -10 (should have been +10) (collimate M1 while slewing)

01:34 pointto M5_0619 ACQUIRE gs=0 PAM- IE=-17 CA=-5 (Note that these last 20 minutes may not have been on the correct star.)

01:36 azslewtohold 372

01:46 pointto M5_0619 ACQUIRE gs=0 PAM- IE=-23 CA=+8

The main result here is that things are thermally moving too much to get a good result on the azimuth scale. It seems to be correct at the 5 arcsec level.

01:57 IE=-26 CA=+14 just to calibrate the drift.

01:59 pointto StoneMagw gs=0 ACQUIRE PAM- (just to see if we can see anything)

02:04 Saw only two faint stars, so try some spiral search, but there are stars everywhere here.

02:10 pointto M5_1469 ACQUIRE PAM- gs=0 (18 +16)

02:15 pointto Arcturus ACQUIRE PAM- gs=0

02:18 ptspiral_irtc.csh Adjust IE=-160 CA=-45 - Saturated even for 1 ms CCD exposure though clouds.

02:29 pointto M5_1469 ACQUIRE PAM- gs=0 - Adjust IE=-170 CA=-49

02:34 pointto StoneMagw gs=0 ACQUIRE PAM- found two apparently dim stars each 8 arsec from the axis.

02:40 Marco sends preset to StoneMagw gs=44. - no significant star found.

02:42 M5_1469 gs=0 ACQUIRE IE=-180 CA=-47

02:45 StoneMagw gs=44 ACTIVE- found star completely at edge of the acquisition field, so there is some pointing problem running around

02:48 StoneMagw gs=0 ACTIVE - same two stars as we found earlier.

02:56 pointto M5_0656 gs=0 ACQUIRE - We wanna check pointing but LFBG Rotator Fails

Fixed Rotator by restarting MCSPU software, but no star visible.

Start Spiral search. Found it ! * IE=-100 CA=-20*

03:31 pointto StoneMagw gs=0 ACQUIRE. Star is there but is a donut.

03:33 try to collimate.

Close

03:36 RH at 90%, T_air = -10.5 degC. We close !

-- JohnHill - 17 Apr 2009
Topic revision: r5 - 18 Apr 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback